What is the diameter of the circle whose area is 36PI
Answer by stanbon(75887)

What is the diameter of the circle whose area is 36PI
Area = (pi)r^2
(pi)r^2 = 36(pi)
r^2 = 36
r = 6
---------
Diameter = 2r = 12 units
===============
Cheers,
Stan H.